OpenStudy (anonymous):

What is the equation of the graphed line? A. y=-5/2x B. y=-2/5x C. y=2/5x D. y=5/2x Please Help , Will Fan And Medal

Vocaloid (vocaloid):

pick two points on the graph and use the slope formula

Vocaloid (vocaloid):

nope

Vocaloid (vocaloid):

(0,0) = (x1,y1) (5,2) = (x2,y2) m = slope = (y2-y1)/(x2-x1) = ?

Vocaloid (vocaloid):

whoops, made a small typo

Vocaloid (vocaloid):



Vocaloid (vocaloid):

not quite

Vocaloid (vocaloid):

y2-y1 = ?

Vocaloid (vocaloid):

 take another look at the diagram